Calculate the monthly payment using this formula ? = P*(1+?)^T/(1+?)^T −1

Principal Loan= 100,000

Interest rate= 3.333% for 30 years

I used the calculate solver and got 439.77 but when I plug it into this equation I get a different number.
Answer #1
PVOrdinary Annuity = C*[(1-(1+i/(f*100))^(-n*f))/(i/(f*100))]
C = Cash flow per period
i = interest rate
n = number of payments I f = frequency of payment
100000= Cash Flow*((1-(1+ 3.333/1200)^(-30*12))/(3.333/1200))
Cash Flow = 439.77
Using Calculator: press buttons "2ND"+"FV" then assign
PV =-100000
I/Y =3.333/12
N =30*12
FV = 0
CPT PMT
Using Excel
=PMT(rate,nper,pv,fv,type)
=PMT(3.333/(12*100),12*30,,100000,)